On January 1, Brothers Corporation issues $10,000,000 of 6%, 15 year bonds, with interest paid semi-annually. The market rate of interest is 8%.

  1. What is the face amount of the bond offering?

10,000,000 * ((1)/(1+.04)^30)) = $3,083,186.68

  1. What will the actual semi-annual interest payment be?

(1-(1+.04)^-30)/(.04) = 17.29 * =

  1. What interest rate is used to compute the present value? Face interest rate or market rate?


  1. How many interest periods will there be in this bond offering?


  1. Based upon the facts provided, will these bonds be sold at a “discount”, “par” or “premium”?
Answer #1

The actual semi-annual interest payment =

Face value * rate of interest * 1/2 =

10000000*6%*1/2= 300000

Market interest rate is used to compute the present value.

Interest periods of bond offering = 15 years * 2 = 30 interest periods

Being the market interest is higher, so the bonds would be sold at discount.

Bonds sale price = semi -int. * PVIFA(4%,30) + Par value * PVIF(4%,30)

Bonds sale price = 300000 * 17.292 + 10000000 * 0.3083 = 8270600